OpenStudy (anonymous):

A triangle has side lengths of 14 cm, 48 cm and 50 cm. Classify it as acute, obtuse, or right

OpenStudy (gudden):

its a right triangle @shangordon using pythagoras it implies since (50)^2 = 48^2 + 14^2 i.e. 2500 = 2304 + 196 and hence its a right angled triangle